### v2

@@ 4.3.3 Ablation study
LVE $\downarrow$ MVE $\downarrow$ FDD $\downarrow$ LVE $\downarrow$ MVE $\downarrow$ FDD $\downarrow$ LVE $\downarrow$ MVE $\downarrow$ FDD $\downarrow$
ST w Wav2Vec2 7.127 1.241 2.349 4.439 2.276 41.374 3.198 2.329 33.671
ST w WavLM 6.773 1.004 2.220 4.333 2.162 31.922 2.907 2.004 8.500
ST w TD 6.512 0.995 1.858 4.827 2.168 36.789 2.136 1.725 35.044
ST w BiGRU 6.584 0.994 2.054 4.421 2.103 41.456 2.534 1.987 34.098
ST w CA Fusion 3.122 0.895 2.137 4.481 2.157 31.125 4.287 2.532 121.00
ScanTalk 6.375 0.987 2.101 4.044 2.057 40.051 2.435 1.678 32.202
5 Conclusions
@@ Acknowledgments
This work is supported by the CNRS International Research Project project GeoGen3DHuman. This work was also partially supported by “Partenariato FAIR (Future Artificial Intelligence Research) - PE00000013, CUP J33C22002830006” funded by NextGenerationEU through the italian MUR within NRRP, project DL-MIG. This work was also partially funded by the ministerial decree n.352 of the 9th April 2022 NextGenerationEU through the italian MUR within NRRP. This work was also partially supported by ANR project 4D SHAPE ANR-24-CE23-5907-01, ANR-24-CE23-5907-03. This work was also partially supported by Fédération de Recherche Mathématique des Hauts-de-France (FMHF, FR2037 du CNRS).

### v3

@@ 4.4 Unregistered Setting
Trained on LVE MVE FDD DTW DFD $\delta_{M}$ $\delta_{Cd}$ LVE MVE FDD DTW DFD $\delta_{M}$ $\delta_{Cd}$ LVE MVE FDD DTW DFD $\delta_{M}$ $\delta_{Cd}$
Rgt (MSE) 6.38 0.99 2.10 1.39 5.71 0.75 0.60 4.04 2.05 40.0 1.47 6.48 0.97 0.60 2.43 1.67 32.20 0.90 4.08 0.23 0.74
Rgt (CH) 6.57 1.01 2.31 1.53 6.24 0.77 0.59 4.30 2.14 34.9 1.54 6.75 0.99 0.62 3.71 2.52 110.6 1.36 5.64 0.35 0.74
Rmsh (CH) 7.13 1.43 8.22 1.49 6.04 0.81 0.71 5.59 2.37 28.2 1.90 8.20 1.24 0.69 5.51 2.86 125.1 1.76 6.41 0.33 0.77
Rmsh+H (CH) 7.13 1.37 9.94 1.73 6.80 0.93 0.71 5.45 2.21 37.1 1.75 7.72 1.04 0.66 5.69 3.00 26.1 1.66 6.43 0.41 0.79
Rmsh+H+GN (CH) 7.91 1.80 39.0 1.75 6.57 0.88 0.68 4.86 2.23 37.2 1.63 7.19 1.04 0.66 5.75 3.11 125.0 1.66 6.44 0.42 0.78
As shown in Table 4, when ScanTalk is trained on unregistered datasets, it still achieves strong performance on registered test data. Although some degradation in results occurs as data quality decreases, the overall metrics remain consistently high, demonstrating the strong generalization capability and robustness of ScanTalk.
Table 5: Distance between generated motions and true scan motions. Here, $\sigma=0.1$ for the varifold metric.
Trained on HD CD ($\times 10^{-5}$) Varifold ($\times 10^{-5}$)
Reg (CH) 0.038 7.60 1.74
Rmsh (CH) 0.042 7.89 1.98
Rmsh+H (CH) 0.044 7.92 2.02
Rmsh+H+GN (CH) 0.039 7.98 2.03
Finally, we evaluate our model performance on raw scan meshes with these different training strategies. We animate the first frame of raw scan sequences from VOCA. We compare the predicted motions against ground truth raw scans sequences. To do this, we use three other unregistered metrics: the mean Chamfer distance, the Hausdorff distance and the varifold distance introduced in Section 3.5.2. As shown in Table 5, training on either registered or unregistered datasets with unregistered losses yields very similar results when tested on raw scans. This further highlights the robustness of ScanTalk across both registered and unregistered settings.
5 Discussions and Conclusions
@@
We acknowledge the ethical implications associated with generating 3D facial animations. The creation of synthetic content using 3D faces carries inherent risks, including the potential for both intentional misuse and unintended societal consequences. We underscore the importance of a human-centered approach in the development and deployment of such technologies. A human-centered design philosophy is critical for ensuring that technology serves and benefits people. The aim of our work is to advance research by addressing an open problem in the literature, with the broader goal of supporting meaningful and beneficial applications. While speech-driven facial animation has many potential uses, some positive, others potentially harmful, we stress the importance of responsible usage and rely on end users to apply this technology ethically and appropriately.
@@ Acknowledgments
This work is supported by the CNRS Int. Research Project GeoGen3DHuman. This work was also partially supported by “Partenariato FAIR (Future Artificial Intelligence Research) - PE00000013, CUP J33C22002830006” funded by NextGenerationEU through the italian MUR within NRRP, project DL-MIG. This work was also partially funded by the ministerial decree n.352 of the 9th April 2022 NextGenerationEU through the italian MUR within NRRP. This work was also partially supported by the project 4DSHAPE ANR-24-CE23-5907 of the French National Research Agency (ANR), and by Fédération de Recherche Mathématique des Hauts-de-France (FMHF, FR2037 du CNRS). This work was also partially supported by the AI4Debunk project (HORIZON-CL4-2023-HUMAN-01-CNECT grant n.101135757).